Write a linear function f with the values f(0)=2 and f(9)=4.
Answer by Boreal(15235)

(0, 2)
(9, 4)
slope is 2/9
point slope y=y1=m(x-x1) m slope (x1, y1) point
y-2=(2/9)x
y=(2/9)x+2